jigglypuff314 (jigglypuff314):

Let v(t) be the velocity, in feet per second, of a skydiver at time t seconds. After her parachute opens, he velocity satisfies the differential equation dv/dt= -2v -32, with initial condition v(0)=-50 Use separation of variables to find an expression for v in terms of t.

jigglypuff314 (jigglypuff314):

Not sure if I did it right, but I got v(t)=-34e^(-2t)-16 ..?

jigglypuff314 (jigglypuff314):

@ganeshie8 @Loser66 could you help?

OpenStudy (sleepyhead314):

@agent0smith @campbell_st @jim_thompson5910 help?

ganeshie8 (ganeshie8):

Looks good !

jigglypuff314 (jigglypuff314):

Ok thanks! :) Then what would I do about lim x->infinity of v(t) ?

ganeshie8 (ganeshie8):

take the limit and see

ganeshie8 (ganeshie8):

v(t)=-34e^(-2t)-16

jigglypuff314 (jigglypuff314):

I forgot how xD

ganeshie8 (ganeshie8):

as "t" increases, the first term reduces right ?

ganeshie8 (ganeshie8):

cuz its a negative exponent

ganeshie8 (ganeshie8):

as \(t \to \infty\), \(e^{-t} = \dfrac{1}{e^t} \to ?\)

jigglypuff314 (jigglypuff314):

Right ... So the -34e^-2t part approaches 0? So I get -16?

ganeshie8 (ganeshie8):

exactly ! the first term vanishes completely, and the velocity approaches -16

jigglypuff314 (jigglypuff314):

Great! :D Thank you! :)

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

if you're hearing it for the first time, its an amazing result actually

ganeshie8 (ganeshie8):

the free falling body cannot continue to increase its velocity forever - the velocity saturates after reaching a certain limit

ganeshie8 (ganeshie8):

google `terminal velocity` for more :)

jigglypuff314 (jigglypuff314):

Makes sense or you could be falling at speed of infinity xD Then "It is safe to land when her speed is 20 ft per sec. At what time does she reach this speed?"

jigglypuff314 (jigglypuff314):

Would that be 20= -34e^-2t -16 ?

ganeshie8 (ganeshie8):

v(t)=-34e^(-2t)-16 -20 =-34e^(-2t)-16 solve t

OpenStudy (sleepyhead314):

yeah but wolfram seems to be confused... http://www.wolframalpha.com/input/?i=20+%3D+-34e%5E%28-2x%29+-16

ganeshie8 (ganeshie8):

the velocity is negative, cuz you're falling down, and the given differential equation assumes that the increasing position is UP.

jigglypuff314 (jigglypuff314):

I'm confused? :/

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

so, velocity is POSITIVE, if u are going in that direction. velocity is NEGATIVE, if u are going opposite to that direction

jigglypuff314 (jigglypuff314):

She's falling up? O.o

ganeshie8 (ganeshie8):

since the given differential equation assumes increasing height is toward UP (why ?), and since she is falling down, that means she is going in opposite direction to the increasing height. so velocity is NEGATIVE.

jigglypuff314 (jigglypuff314):

Erm... So whut do i do?

ganeshie8 (ganeshie8):

set ur solution equal to -20 and solve t

jigglypuff314 (jigglypuff314):

t=1.07 ...?

ganeshie8 (ganeshie8):

thats right ! after 1.07 seconds her speed reaches 20 units/sec

jigglypuff314 (jigglypuff314):

Alright! :) Thank you for your help! <3

ganeshie8 (ganeshie8):

you wlcme :)